pooja195 (pooja195):

Matt and Ming are selling fruit for a school fundraiser. Customers can buy small boxes of oranges and large boxes of oranges. Matt sold 3 small boxes of oranges and 14 large boxes of oranges for a total of $203. Ming sold 11 small boxes of oranges and 11 large boxes of oranges for a total of $220. Find the cost each of one small box of oranges and one large box of oranges.

OpenStudy (igreen):

Matt: 3s + 14l = 203 Ming: 11s + 11l = 220 Where 's' = small boxes and 'l' = large boxes. I think it wants you to solve the two equations..

OpenStudy (igreen):

It's a system of equations: 3s + 14l = 203 11s + 11l = 220

pooja195 (pooja195):

Ah ok i needed help finiding the equations i have it from here thanks :) I have a hard time with finding the equations >_<

OpenStudy (igreen):

Oh, okay.

OpenStudy (anonymous):

:0 oh pooja ;)

OpenStudy (anonymous):

Do you need more help?

pooja195 (pooja195):

nawh i got it from here but welcome to openstudy! :D

OpenStudy (anonymous):

Yeah, welcome to the famous, OPENSTUDY!!!!!!!!!!!!!!!!!!!!!! *oh brother*

OpenStudy (anonymous):

*sighs*

OpenStudy (anonymous):

:)